# thanks: These three authors contributed equallythanks: These three authors contributed equallythanks: These three authors contributed equally

### v1

@@ Appendix E Modeling of time-domain THz data
The reflection spectra obtained by the time-domain technique were modeled using the transfer matrix formalism [54]. Following the experimental configuration used, we considered a plane-parallel slab of MnTe (the thickness of 989 $\mu$m) placed on a highly reflective metallic layer. The dielectric part of the optical response was described as $\epsilon=\epsilon_{bg}[1-\omega_{p}^{2}/(\omega^{2}+i\omega\gamma_{p})]$, where $\omega_{p}=[Ne^{2}/(\epsilon_{0}\epsilon_{\mathrm{bg}}m_{\mathrm{eff}})]^{1/2}$ is the screened plasma frequency. It describes the response due to free carriers (holes) with the density $N$, effective mass $m_{\mathrm{eff}}$, and scattering rate $\gamma_{p}$. To describe the magnon resonance, the magnetic susceptibility was expressed using a Lorentz-type model, $\mu=1+\chi\omega_{m}^{2}/(\omega_{m}^{2}-\omega^{2}-i\omega_{m}\gamma_{m})$, where $\chi$ stands for the oscillator strength of the AFMR, $\omega_{m}$ is the magnon energy and $\gamma_{m}$ is the AFMR line width.
@@
To reduce the number of free-fitting parameters, the effective mass was set to the mass of a bare electron ($m_{\mathrm{eff}}=m_{0}$). Then, we analyzed the temperature dependence of the electrical resistivity (Fig. 3d) to get estimates for $N(T)$ and $\gamma_{p}(T)$. To this end, we fitted the profile of electrical resistivity using an empirical formula, $R(T)=a(1+be^{T/T_{b}})(1+ce^{-T/T_{c}})$. With increasing $T$, this formula describes reasonably well (see the solid line in Fig. 3d) the initial rapid drop in resistivity due to the sharply increasing hole density, which is followed, above $T\approx 50$ K, by a gradual increase in resistivity, driven by the decreasing mobility of the holes. To separate the contributions of the carrier density and the scattering rate to the temperature dependence of (Drude-type) electrical resistivity, $\rho=m_{\mathrm{eff}}\gamma_{p}/(Ne^{2})$, the carrier density at $T=50$ K was anchored at $N=5\times 10^{16}$ cm-3 – the value typically observed in Hall experiments in explored MnTe samples.
In addition to the temperature dependence of the magnon energy, presented and modeled in the main text (Fig. 3c), our fitting procedure also provided us with the width and strength of the AFMR, including their temperature dependencies. The width appeared well reproduced by the effective formula: $\gamma_{m}=\gamma_{m}^{0}(1+T[K]/300)$, where $\gamma_{m}^{0}=40$ $\mu$eV. Regarding oscillator strength, it was highest around 10 K and weakened rapidly with $T$. The magnon excitations became barely observable in the temperature range 50-100 K, but the strength then gradually increased at higher temperatures. This behavior can be traced directly in the bare data set (Fig. 3a) through the splitting of the magnon-polariton modes at the crossing of the AFMR with subsequent modes of the Fabry-Perot modes of the cavity formed by the sample slab. The Rabi splitting is proportional to $\sqrt{\chi}$. The Fabry-Perot pattern also allowed us to firmly determine the background dielectric function that varies smoothly with $T$, $\epsilon_{\mathrm{bg}}(T)=20.5+i0.08+2.0\cdot 10^{-3}(1+i)T[\mathrm{K}]$.
Appendix F High-field high-frequency electron spin resonance - dataset
A complementary experiment was carried out on sample A using the ESR technique. The experimental data collected at $T=5$ K are plotted in Fig. 7. There, we plot the MnTe microwave transmission as a function of the applied magnetic field at three selected microwave frequencies ($\hbar\omega=1.8,2.1$ and 2.7 meV), in the form of $B$-derivatives. Vertical dashed lines mark the positions of the relevant resonances.

### v2

@@ Appendix D Modeling of time-domain THz data
The reflection spectra obtained on sample B by the time-domain technique were modeled using the transfer matrix formalism [58]. Following the experimental configuration, we considered a plane-parallel slab of $\alpha$-MnTe (the thickness of 989 $\mu$m) placed on a highly reflective metallic layer. The dielectric part of the optical response was described as $\epsilon=\epsilon_{bg}[1-\omega_{p}^{2}/(\omega^{2}+i\omega\gamma_{p})]$, where $\omega_{p}=[Ne^{2}/(\epsilon_{0}\epsilon_{\mathrm{bg}}m_{\mathrm{eff}})]^{1/2}$ is the screened plasma frequency. It describes the response due to free carriers (holes) with the density $N$, effective mass $m_{\mathrm{eff}}$, and scattering rate $\gamma_{p}$. To describe the magnon resonance, magnetic susceptibility was expressed using a Lorentz-type model, $\mu=1+\chi\omega_{m}^{2}/(\omega_{m}^{2}-\omega^{2}-i\omega\gamma_{m})$, where $\chi$ stands for the oscillator strength of the AFMR, $\omega_{m}$ is the magnon energy and $\gamma_{m}$ is the width of the AFMR line.
@@
To reduce the number of free parameters in this optical model, we started with the analysis of the temperature dependence of the dc electrical resistivity (Fig. 3d). First, we assumed that this resistivity can be described using a standard Drude-type formula $\rho=m_{\mathrm{eff}}\gamma_{p}/(Ne^{2})$. Second, we used the empirically constructed function: $\rho(T)=a(1+be^{T/T_{b}})(1+de^{-T/T_{d}})$ to reproduce the experimentally observed temperature dependence of electrical resistivity. Good agreement was found for the following parameters: $a=0.39(5)$ $\Omega$m, $b=0.28(5)$, $T_{b}=62(5)$ K, $d=9.1(5)\times 10^{4}$ and $T_{d}=2.4(5)$ K, see the solid line in Fig. 3d. In the range of 10-50 K, this function describes well the rapid drop in resistivity due to the fast increase in the hole density, $N(T)\propto 1/(1+de^{-T/T_{d}})$. On the other hand, the formula reproduces the gradual increase in resistivity above 50 K, driven by the decreasing mobility of the holes, $\gamma_{p}(T)\propto(1+be^{T/T_{b}})$. To separate the values of $N(T)$ and $\gamma_{p}(T)$ in the Drude formula, the carrier density at $T=50$ K was anchored at $5\times 10^{16}$ cm-3 – the value typically observed in Hall experiments in explored samples of $\alpha$-MnTe. Here we assumed that the effective mass is close to that of the bare electron ($m_{\mathrm{eff}}=m_{0}$). At $T=50$ K, this results in $\hbar\omega_{p}=(2\pm 1)$ meV and $\gamma_{p}=(30\pm 15)$ meV. The latter value shows that absorption due to free charge carriers is flat in the THz range, thus creating a smooth background response for a relatively sharp magnon resonance.
@@
In addition to the temperature dependence of the magnon energy, presented and discussed in the main text (Fig. 3c), our modeling procedure also provided us with the width and strength of the AFMR, including their temperature dependence. The width appeared to be well reproduced by the effective formula: $\gamma_{m}=\gamma_{m}^{0}(1+T[K]/100)$, where $\gamma_{m}^{0}=(20\pm 10)$ $\mu$eV. Regarding the oscillator strength, its temperature dependence can be traced directly in the experimental data (Fig. 3a) through the splitting of the magnon-polariton modes. These are formed when the AFMR crosses subsequent modes of the cavity formed by the sample slab itself. These avoided crossings are related to the vacuum Rabi splitting, which is directly proportional to $\sqrt{\chi}$ [59, 60]. In our modeling, only the low-temperature part of the data can be reproduced with a constant oscillator strength. At $T>35$ K, the oscillator strength drops significantly – by an order of magnitude – and partly recovers only above about 100 K. Hence, it seems that the magnon oscillator strength and the dc resistivity have a related temperature dependence. One may even invoke the possibility that the magnon oscillator strength is related to the density of thermally excited holes. However, further discussion goes beyond the scope of this paper. For our modeling, we used a heuristic function $\chi(T)$: a constant low-temperature value of $6\times 10^{-4}$ that continuously changes into a $\chi(T)\propto R(T)/T^{3}$ profile in the range of $T=35-300$  K. The Fabry-Perot pattern also allowed us to establish the background dielectric function that varies smoothly with $T$, $\epsilon_{\mathrm{bg}}(T)=20.6+i0.1+8\cdot 10^{-4}(1+i)T[\mathrm{K}]$.
Appendix E High-field high-frequency electron spin resonance - dataset
A complementary experiment was carried out on sample A using the ESR technique. The experimental data collected at $T=5$ K are plotted in Fig. 8. There, we plot the $\alpha$-MnTe microwave transmission as a function of the applied magnetic field at three selected microwave frequencies ($\hbar\omega=1.8,2.1$ and 2.7 meV), in the form of $B$-derivatives. Vertical dashed lines mark the positions of the relevant resonances.